OpenStudy (anonymous):

[Part 1 (4 points)] Show all work in dividing (4t^2-1/15t+10)/(4t^2-4t+1/3t^2+5t+2)

OpenStudy (anonymous):

\[\frac{4t^2-1}{15t+10}\times \frac{3t^2+5t+2}{4t^2-4t+1}\] is a start then factor as much as possible \[\frac{(2t+1)(2t-1)(t+1)(3t+2)}{5(3t+2)(2t-1)(2t-1)}\] and then cancel away i have seen this exact problem before where does it come from?